The area of a rectangle of length x is given by 3x^2+5x. Find the width of the rectangle.
Answer by ptaylor(1894)

Area(A) of a rectangle equals Length(L) times Width(W) or A=L*W
Now we are told that:
A=3x^2+5x=L*W=x*W divide both sides by x
W=3x+5
CK
L*W=x(3x+5)=3x^2+5x
Hope this helps---ptaylor